# Adaptive Electronic Skin Sensitivity for Safe Human-Robot Interaction

### v1

@@ I Introduction
Although the sense of touch has been somewhat underestimated and researched much less compared to vision, for example, many disciplines have now come to realize its enormous importance for human cognition (e.g., [1]). Touch and the related sense of pain is also critical for protection of our bodies against harmful contacts and impacts. Moreover, “humans are born as tactile creatures. Physical touch is one of the most basic forms of human communication.” [2] With the recent advent of artificial skin technologies covering large areas of their bodies, it is time to endow robots with the sense of touch and leverage it for physical and social Human-Robot Interaction (HRI).
The importance of endowing robots with touch has been recognized for several decades and a large number of technologies have been developed. The focus has largely been on tactile sensing for manipulation, as equipping robot hands/fingers only with tactile sensors requires only relatively small patches of electronic skin. Whole-body artificial skins for robots have been less common, with only few successful solutions such as [3, 4] or those coming from the ROBOSKIN project [5] and deployed most notably on the iCub humanoid robot [6]. The potential of sensing over the whole body surface has also not been fully explored, but Cheng et al. provide an overview of applications of their electronic skin [7]. Artificial skin solutions are also starting to find their way to the collaborative robot industry through, for example, AIRSKIN (detecting contact) or Bosch APAS (using proximity).
@@
This work is mainly motivated by safe HRI, namely how to set the skin sensitivity thresholds such that safety is warranted as per the Power and Force Limiting regime of HRI [9] but maximizing the robot performance (speed). Specifically, we worked with the UR10e robot equipped with the AIRSKIN robotic skin—see Fig. 1. By examining collision during a robotic task in different axis of motion, we show that it is beneficial to set different thresholds for different parts of the robot, and the most promising way is to set these thresholds adaptively based on current velocity and joint configuration of the robot.
The simulation, code and recorded experiment data are publicly available at https://github.com/ctu-vras/adaptive-skin and https://osf.io/3adrm/.
@@ II Related work
This Section provides a brief overview of selected works in Physical Human-Robot Interaction (pHRI). We also present state of the art on the topic of pain in robotics and some literature on artificial skin development, with an emphasis on pain.
@@
The state-of-the-art research in this field goes in multiple directions. One possibility is to combine other pHRI modes, such as Speed and Separation Monitoring (SSM), with PFL to obtain more general approaches [15, 16]. A different approach lies in using additional sensors. Xu et al. [17] proposed an active pipeline that is able to reduce impact forces using the visual-tactile approach. Similarly, Svarny et al. performed an extensive set of experiments to show the impact of robotic skin on collision forces [18].
Robotic skins are generally a promising asset in the future of human-robot interaction. With an increase in the number of available whole-body robotic skins [7], new approaches can be invented to control robots or handle collisions. For example, Guadarrama-Olvera et al. [19] proposed an approach for a pressure-driven controller of a humanoid robot based on the external force and also the size of the contact. An interesting subcategory are neuro-inspired robotic skins. An example of this category is the skin by Liu et al. [20] which consists of synaptic transistors. The authors demonstrated that the skin is able to acquire pain-like reflexes. Moreover, human skin also has the ability to sense thermal stimuli. Neto et al. [21] developed thermoreceptors-based skin that is able to learn at the hardware level to respond to temperature impulses. We refer the readers to the review [22] for more details on bio-inspired robotic sensing.
@@
Our work lies at the intersection of the fields mentioned above. We aim to use standard, pressure-based whole-body robotic skin to introduce the notion of onset of pain in robotics and empirically evaluate its importance and advantages to pHRI.
@@ III-A Robot and Skin
We use the UR10e 6-DOF robotic arm. The robot is a collaborative one, i.e., it can detect collisions and other distortions by itself. Thus, we selected the less restrictive setting of its safety system, so all collisions are firstly detected by our program and skin, and not the robot. We control the real robot through its official Robot Operating System (ROS) Noetic driver in velocity mode, i.e., we send velocities directly to individual joints of the robot.
The robot is further equipped with a robotic skin that we use for our evaluation. Specifically, we use AIRSKIN111https://www.airskin.io/ robotic skin. AIRSKIN detects changes in pressure inside each of its parts—we call them pads. Our robot with 11 pads can be seen in Fig. 1. By default, the system should be connected directly to the robot safety controller, and only binary information (collision / no collision) is available. Thanks to the help of AIRSKIN manufacturers, we can connect the skin to our computer and read the pressure values of the individual pads with the frequency of . The skin readout is also integrated to ROS.
@@ III-B Simulation, Software and Data
On top of the real robot, we also created a simulation for both the robot and the skin. The simulation uses PyBullet [25] as a background physics engine. We created a front-end interface that contains the UR10e robot with simulated AIRSKIN. It can work in two modes:
native mode, where user has only low-level control of the robot without any external planners or controllers. The simulation can run off-screen or with a custom visualizer created with Open3D Python library [26]. In this mode, the simulation speed is in orders of magnitude faster than real-time (depending on the hardware) and can be used, for example, for reinforcement learning;
@@ III-C Performed Task
We decided to simulate a pick-and-place task, whereby the robot motion are disturbed by collisions, simulating physical human-robot collaboration. We designed the task in such a fashion that the robot first moves in its y-axis (forward from the base), then in the negative and positive z-axis (up and down), and finally in the negative x-axis (sideways)—see Fig. 2. The task has been designed in this way, as it can be both theoretically and empirically shown that the collision forces will depend not only on the link speed but also on the collision direction and position in the workspace (see e.g., [27, 18]). Furthermore, all joints of the robot are employed during the task.
In every moment, we assume that every contact is a transient one. The ISO/TS 15066 [9] described this as a “dynamic impact”, where the human body part impacted by the robot is not constrained and can retract from the robot, resulting in a shorter impact duration.
@@ III-D Collision Reaction
III-E Onset of Pain
We discretized the “robot pain” to three thresholds.
Using these thresholds, we tested four settings of the onset of pain on individual skin parts (pads): (1) the same threshold on every pad (UNIFORM); (2) different thresholds on pads (BODY PARTS); (3) adaptive threshold based on ISO/TS 15066 (LINK VELOCITY); (4) adaptive threshold based on ISO/TS 15066 and effective mass of the robot (EFFECTIVE MASS).
Settings (1) and (2) are static, i.e., the same threshold is set during the whole task. Settings (3) and (4) are dynamic and the thresholds for individual skin pads are set with the frequency of (every ).
@@ III-E1 UNIFORM
In this setting, there is no onset of pain, since all skin pads have the same threshold for pain, specifically the most sensitive. We consider this metric as a baseline as it is a common setting.
@@ III-E2 BODY PARTS
We took inspiration from humans and ISO/TS 15066, and divided the robot into three parts, similar to the human arm—upper and lower arm, and hand. This division is not only similar to humans, but also intuitive when looking at the robot, as it has two longer sections and then the gripper—see Fig. 1. We examined velocities of the given body parts during the task and observed that the hand and lower arm usually move with higher velocity than the upper arm. Therefore, we set the thresholds for this mode as the most sensitive for hand and lower arm, and the middle threshold for upper arm.
@@ III-E3 LINK VELOCITY
This settings of pad onsets is based on the equation for maximal velocity from ISO/TS 15066. The equation determines the maximal velocity using mass-spring-mass model as
@@
We use Eq. 4 to compute the possible force exerted when colliding with each skin pad whenever a new joint state of the robot is available. Again, based on ISO/TS 15066 we set (spring constant of the back of non-dominant hand) and to the half of the moving mass of the robot up to the given skin part. Then, based on maximal allowed force for transient contact () we selected the threshold of pain for the given pad as:
@@ III-E4 EFFECTIVE MASS
This section uses a term called effective mass. This property relates current configuration and dynamics of the robot. Based on the work of Khatib [30], one can compute the effective mass in the direction as
@@ IV Experiments and Results
This Section summarizes the experiments done. We conducted the same set of experiments in two setups, the real world and the simulation. We tested two scenarios—stopping (STOP) after contact and moving away from the contact (AVOID). We saved all data from the robot (including skin pressures) in the form of rosbags—available at https://osf.io/3adrm/.
@@ IV-A Experiment Description
In both environments, we assume transient contact with a human part with a weight of (weight in Eq. 4) and we set the Cartesian velocity of the end-effector link to . This velocity was selected as the borderline velocity that in some configurations exceeds the limit of ISO/TS 15066 when computed using Eq. 4.
During the experiments, the robot always performs its task described in Section III-C and we let it collide with an obstacle. In each run, only one collision can happen. We examine collision during movement in each direction of the task (+y, -z, and -x axis) and on each of the body parts of the robot (upper arm, lower arm, and hand). As there are several skin pads on each of the body parts, we always select one pad that moves in the most similar direction compared to the end-effector of the robot. For each threshold setting (see Section III-E) there are 9 individual runs (1 per each axis; 3 per body part), making 36 runs in total per reaction behavior (STOP and AVOID). We performed 10 repetitions of each experiment in the simulation and 3 repetitions in the real world, resulting in 720 and 216 runs.
@@ IV-B Simulation Experiments
Arguably, the most important property is the average run-time of the experiments based on the thresholds used. We provide bar graphs in Fig. 3 and numerical values in Table I. As described in the beginning of this section, the maximum stopping time of the robot is . We assume that it would not be the case in real collaborative task, as there would probably be a prolonged period of time when the robot is stopped, because someone has to re-enable its movement. Therefore, we add to each run when the robot detects a collision. The reason we did not do it directly in the simulation is a pure endeavor to make the experiments faster.
@@
In case of AVOID, there is a small difference even between UNIFORM and BODY PARTS (about ), which is probably caused by the fact that because the reaction time is higher, the direction of movement of the given link can be slightly different. It may not seem as a significant influence on total time, but because the only body part with different pain threshold is the upper arm, the avoid movement is constrained only to the first two joints of the robot. Thus, even a small difference in movement direction can change avoidance velocity. The rest is analogous to the STOP scenario.
For both reaction strategies, the SDs are higher for dynamic threshold switching strategies. It is an instinctive result, as these two have a higher deviation of stopped and non-stopped runs.
Another property we can examine is the reaction time to collision. The values are shown in Fig. 4. The results correspond to our hypothesis, that is, they should be the lowest for UNIFORM and the highest for EFFECTIVE MASS. The reason is that with UNIFORM, all thresholds are set to the most strict ones, thus the robot should stop faster. On the other hand, with LINK VELOCITY and EFFECTIVE MASS it is expected that there will be less strict thresholds in most cases, naturally resulting in higher reaction times. We can confirm this assumption using Table III or Fig. 1, where we can see that for LINK VELOCITY and EFFECTIVE MASS, threshold 0 is set only in 35% and 0% of cases, respectively. The overall highest reaction is in LINK VELOCITY mode. It may seem high at first glance, but the correlation coefficient between reaction times and the force estimated using Eq. 4 is -0.86. It shows a high negative correlation, i.e., that estimated force is higher for low reaction times and vice-versa. In other words, when the reaction time is high, the estimated force is low. The average estimated force for cases where the reaction time is higher than is , which is far below the limit of . However, it is important to take this into account, as it may cause a problem in some cases.
The last property we explore is the effect of thresholds on distance needed to go away from the collision—naturally, this is only evaluated in AVOID. The results are in Fig. 5. We can see that, similarly to the total time, the traveled distance is the highest for UNIFORM and the lowest for EFFECTIVE MASS. Because LINK VELOCITY and EFFECTIVE MASS assign higher pain thresholds (see Table III and Fig. 1), the robot needs to travel a shorter distance to get away from the collision. In our case, the robot immediately started to come back to trajectory and continue the task, and thus the hypothetically impacted human would have less time to move because of the shorter distance. However, we assume that in the real world there would be, again, some time when the robot would remain stopped and then a shorter avoidance distance means less disruption of tasks.
@@ IV-C Real Setup
In the real world, the impacts were recorded with real humans. To assign pain (pressure) values to individual thresholds, we recorded several values with CBSF-75-Basic, which is a certified measuring device for validation of collaborative tasks with the range . To simulate transient contact, we furthermore used a construction consisting of moving mass and bearing sliders—Fig. 6(a). The mass of the whole moving mass (used for Eq. 4) is . We kindly refer the reader to [18] for more details on the construction. Using this setup, we measured impact forces in x- and y-axis movements of our experiments in hand and lower body regions of the robot. We cannot measure forces when moving in the negative z-axis, as gravity would pull the moving mass to the ground, and we were unable to measure forces on upper arm because of limited space not allowing to properly hit the measuring device. Based on the measurement, we selected the most sensitive threshold as a value that always reacted to the collision with the device; the second threshold as the last value that reacted and the least sensitive threshold circa as 90% of the pressure range of the skin.
@@ V Conclusion, Discussion, and Future Work
We conducted an extensive set of experiments, both in simulation (that we created) and in the real world, to find out how thresholds for pain in robots with whole-body skin should be set. The experiments showed that the pain threshold influences mainly the sensitivity to contacts and collision detection reaction time. These two indirectly influence task properties like total task time or distance needed to travel away from collision. In the beginning, we asked two questions: Should there be different sensitivity (“onsets of pain”) thresholds for different parts of the robot skin? Based on what criterion should the onsets be set? The answer to the first one is straightforward—yes, the onsets of pain should be different. Individual links of the robot move with different velocity and have different weight, therefore they have capability to exert a different force during impact. The visualization of changes in thresholds over time during the task can be seen in Fig. 1. Using different thresholds helps mainly to keep the task interrupted and smooth in case of small, harmless collisions.

### v2

@@ I Introduction
The importance of endowing robots with touch has been recognized for several decades and a large number of technologies have been developed. The focus has largely been on tactile sensing for manipulation, as equipping robot hands/fingers only with tactile sensors requires only relatively small patches of electronic skin. Whole-body artificial skins for robots have been less common, with only few successful solutions such as [1, 2, 3] or those coming from the ROBOSKIN project [4] and deployed most notably on the iCub humanoid robot [5]. A more recent example is the Punyo-1 robot [6]. The potential of sensing over the whole body surface has also not been fully explored, but Cheng et al. provide an overview of applications of their electronic skin [7]. Artificial skin solutions are also starting to find their way to the collaborative robot industry through, for example, AIRSKIN (detecting contact) or Bosch APAS (using proximity).
@@
This work is mainly motivated by safe Human-Robot Interaction (HRI), namely how to set the skin sensitivity thresholds such that safety is warranted as per the Power and Force Limiting regime (PFL) of HRI [9] but maximizing the robot performance (speed). Specifically, we worked with the UR10e robot equipped with the AIRSKIN robotic skin—see Fig. 1. Collision characteristics (contact force, duration, etc.) depend on the speed of the colliding robot link but also on its effective mass, which is a function of the whole robot configuration and the impact direction—the robot is more “massy” in certain configurations and collision directions. Thus, to warrant contacts complying with the PFL limits but at the same time maximizing productivity, collision thresholds that will trigger robot stop or avoidance should be set individually for different parts of the robot bodies and dynamically on the run.
The simulation, code and recorded experiment data are publicly available at https://github.com/ctu-vras/adaptive-skin and https://osf.io/3adrm/. The video is available at https://youtu.be/kWt2J1Qgzpk.
@@ II Related work
This Section provides a brief overview of selected works in Physical Human-Robot Interaction (pHRI), protective electronic skins, and “artificial sense of pain”.
@@
The PFL regime can be also combined with other pHRI modes, such as Speed and Separation Monitoring (SSM), further boosting productivity [15, 16]. A different approach lies in using additional sensors. Xu et al. [17] proposed an active pipeline that is able to reduce impact forces using a visual-tactile approach.
Robotic skins are generally a promising asset in the future of human-robot interaction. With an increase in the number of available whole-body robotic skins [7], new approaches can be invented to control robots or handle collisions. Similarly, Svarny et al. performed an extensive set of experiments to show the impact of robotic skin on collision forces [18]. Guadarrama-Olvera et al. [19] proposed an approach for a pressure-driven controller of a humanoid robot based on the external force and also the size of the contact. An interesting subcategory are neuro-inspired robotic skins. An example of this category is the skin by Liu et al. [20] which consists of synaptic transistors. The authors demonstrated that the skin is able to acquire pain-like reflexes. Moreover, human skin also has the ability to sense thermal stimuli. Neto et al. [21] developed thermoreceptors-based skin that is able to learn at the hardware level to respond to temperature impulses. We refer the readers to the review [22] for more details on bio-inspired robotic sensing.
@@
This work demonstrates how adaptive collision sensitivity thresholds can bring safety and productivity under one hood.
@@ III-A Robot and Skin
We use the UR10e 6-DOF robotic arm. The robot is a collaborative one, i.e., it can detect collisions and other distortions by itself. We selected the least restrictive setting of its safety system, so that all collisions are firstly detected by the sensitive skin and our program and skin and not the robot controller from the manufacturer. We control the real robot through its official Robot Operating System (ROS) Noetic driver in velocity mode, i.e., we send velocities directly to individual joints of the robot.
The robot is also equipped with a robotic skin AIRSKIN111https://www.airskin.io/. AIRSKIN detects changes in pressure inside each of its parts—we call them pads. Our robot with 11 pads can be seen in Fig. 1. By default, the system should be connected directly to the robot safety controller, and only binary information (collision / no collision) is available. With a special airskin maintenance controller, we can connect the skin to our computer and read the pressure values of the individual pads with the frequency of . The skin readout is also integrated to ROS. The standard operation of AIRSKIN, which triggers robot stop after contact is detected through the dedicated safety lines, was not used here.
@@ III-B Simulation, Software and Data
Next to the real robot setup, we also created a simulation for both the robot and the skin. The simulation uses PyBullet [25] as a background physics engine. We created a front-end interface that contains the UR10e robot with simulated AIRSKIN. It can work in two modes:
native mode, where user has only low-level control of the robot without any external planners or controllers. The simulation can run off-screen or with a custom visualizer created with Open3D Python library [26]. In this mode, the simulation speed is orders of magnitude faster than real-time (depending on the hardware) and can be used, for example, for reinforcement learning;
@@ III-C Performed Task
We simulated a pick-and-place task, whereby the robot motion is disturbed by collisions, simulating physical human-robot collaboration. The robot first moves in its y-axis (forward from the base), then in the negative and positive z-axis (up and down), and finally in the negative x-axis (sideways)—see Fig. 2 and the video at https://youtu.be/kWt2J1Qgzpk. The task has been designed in this way, as it can be both theoretically and empirically shown that the collision forces will depend not only on the link speed but also on the collision direction and position in the workspace (see e.g., [27, 18]). All joints of the robot are employed during the task.
In every moment, we assume that every potential contact is a transient one. The ISO/TS 15066 [9] described this as a “dynamic impact”, where the human body part impacted by the robot is not constrained and can retract from the robot, resulting in a shorter impact duration.
@@ III-D Collision Reaction
III-E Collision Sensitivity Thresholds
We considered three collision sensitivity thresholds and four different ways of distributing them on the skin parts (pads) on the robot body: (1) the same threshold on every pad (UNIFORM); (2) different thresholds on pads (BODY PARTS); (3) adaptive threshold based on ISO/TS 15066 (LINK VELOCITY); (4) adaptive threshold based on ISO/TS 15066 and effective mass of the robot (EFFECTIVE MASS).
Settings (1) and (2) are static, i.e., the same threshold is set during the whole task. Settings (3) and (4) are dynamic and the thresholds for individual skin pads are set with the frequency of (every ). The choice of the frequency is affected by the time needed to compute the dynamic properties of the robot.
@@ III-E1 UNIFORM
In this setting, all skin pads have the same threshold, specifically the most sensitive. We consider this metric as a baseline as it is a common setting when protective sensitive skins are employed.
@@ III-E2 BODY PARTS
We divided the robot into three parts, similar to the human arm—upper arm, lower arm (forearm), and hand. This division is not only similar to humans, but also intuitive when looking at the robot, as it has two longer sections and then the gripper—see Fig. 1. We examined the velocities of the given body parts during the task and observed that the hand and lower arm usually move with a higher velocity than the upper arm. Therefore, we set the thresholds for this mode as the most sensitive for the hand and lower arm, and the middle threshold for the upper arm.
@@ III-E3 LINK VELOCITY
This settings of pad onsets is based on the equation for maximal velocity from ISO/TS 15066. The equation determines the maximal velocity using a mass-spring-mass model as
@@
We use Eq. 4 to compute the possible force exerted when colliding with each skin pad whenever a new state of the robot is available (and hence every link will have a new velocity). Again, based on ISO/TS 15066 we set (spring constant of the back of non-dominant hand; the most sensitive human body part according to [9], except for the head) and to the half of the moving mass of the robot up to the given skin part. Then, based on the maximal allowed force for transient contact () we selected the sensitivity thresholds for the given pad as:
@@ III-E4 EFFECTIVE MASS
This property relates the current configuration and dynamics of the robot. Based on the work of Khatib [30], one can compute the effective mass in the direction as
@@ IV Experiments and Results
We conducted the same set of experiments in two setups—in the real world and in simulation—with the four sensitivity threshold settings. We tested two scenarios—stopping (STOP) after contact and moving away from the contact (AVOID). We saved all data from the robot (including skin pressures) in the form of rosbags—available at https://osf.io/3adrm/.
@@ IV-A Experiment Description
In both environments, we assume transient contact with a human body part with a weight of (corrresponding to the human arm; weight in Eq. 4) and we set the Cartesian velocity of the end-effector link to . This velocity was selected as the borderline velocity that in some configurations exceeds the limit of ISO/TS 15066 when computed using Eq. 4.
During the experiments, the robot always performs its task described in Section III-C and we let it collide with an obstacle. In each run, only one collision can happen. We examine collision during movement in each direction of the task (+y, -z, and -x axis) and on each of the body parts of the robot (upper arm, lower arm, and hand). As there are several skin pads on each of the body parts, we always select one pad that moves in the most similar direction compared to the end-effector of the robot. For each threshold setting (see Section III-E) there are 9 individual runs (1 per each axis; 3 per body part), making 36 runs in total per reaction behavior (STOP and AVOID). We performed 10 repetitions of each experiment in the simulation and 3 repetitions in the real world, resulting in 720 and 216 runs, respectively.
@@ IV-B Simulation Experiments
The most important performance measure is the average run-time of the experiments based on the thresholds used. We provide bar graphs in Fig. 3 and numerical values in Table I. In the simulated experiments, when a collision exceeding the sensitivity threshold was detected, the robot stopped for . However, in a real collaborative task, such events typically result in longer task interruptions, because an operator needs to re-enable the robot motion. Therefore, for performance evaluation, we add to each run when the robot detects a collision.
@@
In case of AVOID, there is a small difference even between UNIFORM and BODY PARTS (about ), which is probably caused by the fact that because the reaction time is higher, the direction of movement of the given link can be slightly different. It may not seem as a significant influence on total time, but because the only body part with different sensitivity threshold is the upper arm, the avoid movement is constrained only to the first two joints of the robot. Thus, even a small difference in movement direction can change avoidance velocity. The rest is analogous to the STOP scenario.
For both reaction strategies, the SDs are higher for dynamic threshold switching strategies. This would be expected, as these two have a higher deviation of stopped and uninterrupted runs.
Another property we can examine is the reaction time to collision. The values are shown in Fig. 4. The results correspond to our hypothesis, that is, they should be the lowest for UNIFORM and the highest for EFFECTIVE MASS. The reason is that with UNIFORM, all thresholds are set to the most strict ones, thus the robot should stop faster. On the other hand, with LINK VELOCITY and EFFECTIVE MASS it is expected that there will be less strict thresholds in most cases, naturally resulting in higher reaction times. We can confirm this assumption using Table III or Fig. 1, where we can see that for LINK VELOCITY and EFFECTIVE MASS, threshold 0 is set only in 35% and 0% of cases, respectively. The overall longest reaction time is in LINK VELOCITY mode. It may seem high at first glance, but the correlation coefficient between reaction times and the force estimated using Eq. 4 is -0.86. It shows a high negative correlation, that is, that the estimated force is higher for the short reaction times and vice versa. I The average estimated force for cases where the reaction time is higher than is , which is far below the limit of .
The last property we explore is the effect of thresholds on distance needed to go away from the collision—naturally, this is only evaluated in AVOID. The results are in Fig. 5. We can see that, similarly to the total time, the traveled distance is the highest for UNIFORM and the lowest for EFFECTIVE MASS. Because LINK VELOCITY and EFFECTIVE MASS assign higher sensitivity thresholds (see Table III and Fig. 1), the robot needs to travel a shorter distance to get away from the collision. In our case, the robot immediately started to come back to trajectory and continue the task, and thus the hypothetically impacted human would have less time to move because of the shorter distance. However, we assume that in the real world there would be, again, some time when the robot would remain stopped and then a shorter avoidance distance means fewer interruptions of the task.
@@ IV-C Real Setup
In the real world, the impacts were recorded with a human experimenter. To find AIRSKIN pressure values approximately corresponding to the force thresholds, we recorded several values with CBSF-75-Basic, which is a certified measuring device for validation of collaborative tasks with the range . To simulate transient contact, we furthermore used a construction consisting of moving mass and bearing sliders—Fig. 6(a). The mass of the whole moving mass (used for Eq. 4) is . We kindly refer the reader to [18] for more details on the construction. Using this setup, we measured impact forces in x- and y-axis movements of our experiments in hand and lower body regions of the robot. We cannot measure forces when moving in the negative z-axis, as gravity would pull the moving mass to the ground, and we were unable to measure forces on upper arm because of limited space not allowing to properly hit the measuring device. Based on the measurement, we selected the most sensitive threshold as a value that always reacted to the collision with the device; the second threshold as the last value that reacted and the least sensitive threshold circa as 90% of the pressure range of the skin.
@@ V Conclusion, Discussion, and Future Work
We conducted an extensive set of experiments in simulation, which we created, and in the real world, to find out how collision sensitivity thresholds robots should be set in order to warrant safety and maximize productivity. The experiments showed that the sensitivity thresholds determine the proportion of contacts that do not interrupt robot task execution and also collision detection reaction time. These two indirectly influence task properties like total task time or distance needed to travel away from collision. In the beginning, we asked two questions: Should there be different sensitivity (“onsets of pain”) thresholds for different parts of the robot skin? Based on what criterion should the onsets be set? The answer to the first one is straightforward—yes, the collision sensitivity thresholds should be different. Individual links of the robot move with different velocity and have different weight, therefore they have capability to exert a different force during impact. The visualization of changes in thresholds over time during the task can be seen in Fig. 1. Using different thresholds helps mainly to keep the task uninterrupted and smooth in case of small, harmless collisions.